Opened 9 years ago

Closed 8 years ago

Last modified 9 months ago

#372 closed flight processing (fixed)

EUFAR10/01 (AIMWETLAB), flight day 235b/2010, Lake Balaton Peninsula

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by iopa)

Data Location: ~arsf/workspace/EUFAR10_01-2010_235b_Lake_Balaton_Peninsula

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of species composition and stand health of shore vegetation, mapping of shore elevation and inundation structure.

This flight will also be used for EUFAR10-07.
Scientific Objective: High accuracy land use/land cover mapping, forest type mapping and invasive shrub species location

Priority: a4h

PI: A.Zlinszky

Note; the symlinks from the workspace point to the archive-overflow in ~arsf/arsf_data/2010/flight_data/ and therefore will break when space is available for the project to be moved out of there.

Sensors:

  • Eagle (28/03/2011)
  • Hawk (28/03/2011)
  • Leica LIDAR (04/03/2011)
  • Digital Photography (28/03/2011)

Attachments (1)

2352010.sync (330 bytes) - added by iopa 9 years ago.
Eagle & Hawk SCT offsets

Download all attachments as: .zip

Change History (47)

comment:1 Changed 9 years ago by iopa

Please note that eagle files 5-7 still need to be downloaded from thelma.

comment:2 Changed 9 years ago by knpa

  • Description modified (diff)

Added scientific objective.

comment:3 Changed 9 years ago by knpa

  • Description modified (diff)

This flight will also be used for EUFAR10-07.

comment:4 Changed 9 years ago by iopa

From an email by Phil Goy on 19/10/2010,
Eagle files VNIR235b-10-5.raw,VNIR235b-10-6.raw,VNIR235b-10-7.raw
are unrecoverable and we will have to process only the partial
ones we already have.

The headers for those still need to be fixed to match the partial raws.

comment:5 Changed 9 years ago by iopa

The headers have been fixed and the lines are processable as any other, BUT

for lines 5-7;
azspec_args_extra needs to contain (at least):
azspec_args-extra = -sdr eagle/darkFrames.bil -sdh eagle/darkFrames.bil.hdr

Those dark frames have been extracted from VNIR235b-10-1.raw, it has been tested as far as azspec and seems to work. If any problems arise in next stages for those lines, keep this in mind.

comment:6 Changed 9 years ago by iopa

The above is for eagle only.

comment:7 Changed 9 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:8 Changed 9 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking camera.

comment:9 Changed 9 years ago by knpa

All images look OK, no major over/under exposure or shots of bay doors.

comment:10 Changed 9 years ago by knpa

Doing ipas.

comment:11 Changed 9 years ago by knpa

Completed ipas using lidar lever arm.

Still some dispute over correct camera lever arm, will do camera version of sol file when this has been resolved.

Taking Lidar.

comment:12 Changed 9 years ago by knpa

Lidar complete and ready for delivery check at:

~airborne/workspace/EUFAR10_01-2010_235b_Lake_Balaton_Peninsula/delivery/20110211/

comment:13 Changed 9 years ago by iopa

Checking LiDAR.

comment:14 Changed 9 years ago by iopa

LiDAR check;

delivery is in order and ready to go.

As mentioned in the readme, some noise remains unclassified over water, which is outside the area of interest.

comment:15 Changed 9 years ago by knpa

  • Description modified (diff)

Fixed PI name misspelling.

comment:16 Changed 9 years ago by anch

  • Description modified (diff)

Camera, and hyperspectral sent to PI

comment:17 Changed 9 years ago by anch

  • Description modified (diff)

Ignore the above comment.

Actually is was the LiDAR sent to PI.

Hyperspectral and Camera are still to be delivered.

comment:18 Changed 9 years ago by iopa

Taking hyperspectral.

comment:19 Changed 9 years ago by knpa

Camera delivery has been created but there are some photos not over a flightline that have been missed and there is a flightline included in kml that shouldn't be (it's a void line).

comment:20 Changed 9 years ago by knpa

Fixed above points.

Camera ready for delivery check at: /users/rsg/airborne/workspace/EUFAR10_01-2010_235b_Lake_Balaton_Peninsula/delivery/20110304/EUFAR10_01

comment:21 Changed 9 years ago by anch

Delivery check complete.

All fine and ready to go

comment:22 Changed 9 years ago by iopa

Hyperspectral ready for checking at ;

~airborne/workspace/EUFAR10_01-2010_235b_Lake_Balaton_Peninsula/delivery/20110314/EUFAR10_01

Changed 9 years ago by iopa

Eagle & Hawk SCT offsets

comment:23 Changed 9 years ago by anch

  • Summary changed from EUFAR10/01, flight day 235b/2010, Lake Balaton Peninsula to EUFAR10/01 (AIMWETLAB), flight day 235b/2010, Lake Balaton Peninsula

comment:24 Changed 9 years ago by emca

Delivery checked hyperspectral data. All looks good, ready to deliver.

comment:25 Changed 9 years ago by iopa

  • Description modified (diff)

Hyperspectral & Photography have been dispatched (28/03/2011)

comment:26 Changed 9 years ago by iopa

  • Description modified (diff)

comment:27 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving

Project r-synced, ready to be archived.

comment:28 Changed 9 years ago by knpa

Beginning archiving.

Some applanix/gps files are missing.

comment:29 Changed 9 years ago by knpa

Tabi data present.

GRIMM data present

comment:30 Changed 8 years ago by knpa

Tarball created, made available to NEODC (17/05/11)

comment:31 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion

Archived and backed up by NEODC.
Moved to archived disk.

comment:32 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:33 Changed 8 years ago by knpa

Starting reprocessing hyperspectral with APL.

comment:34 Changed 8 years ago by knpa

For some reason the date in the Eagle 13 raw header had been changed to 22nd august rather than 23rd like all the others. As there is no reason in the ticket for doing this and it appears to be incorrect, I am changing it back.

comment:35 Changed 8 years ago by knpa

Ran stage 2.

Slight trickyness to this project as the first line (e+h) is split into two.

comment:36 Changed 8 years ago by knpa

Re-ran stage 2 as apl chain has been updated to generate additional masked files.

comment:37 Changed 8 years ago by knpa

Re-ran Hawk due to additional changes to aplcal in relation to the bad pixel file.

Also re-ran Eagle 1 (split line), eagle 5 (no sync), eagle 6 (dark frame issue), eagle 7 (dark frame issue).

comment:38 Changed 8 years ago by knpa

Not sure if can correct sync for e5 as this line is over water.

Reprocessing everything using the UTM33 datum instead of 34 as this was used previously (and for most other projects) and will allow the different datasets to line up in e.g. gtviewer.

comment:39 Changed 8 years ago by knpa

Managed to correct e5 using water distortions.
Also corrected h5.

e5: 1.5
h5: 1.54

comment:40 Changed 7 years ago by emca

Delivery checked reprocessed hyperspectral data.

Only issues are DEM name (-instead of _) to mention Eagle dropped scans in Readme.

Otherwise all is good.

comment:41 Changed 7 years ago by emca

Changed DEM name and updated Readme, mapped files have been zipped

Reprocessing complete - adding delivery to data to be sent to Zlinszky

comment:42 Changed 7 years ago by emca

Data sent to A Zlinzky 16/08/2012

comment:43 Changed 9 months ago by wja

Archiving
Camera delivery JSONs are failing to be created due to photo filenames in eventfile having quotation marks (").
A new camera delivery has been corrected with this issue amended.
All JSON have been produced and added to PostGIS. amended.

comment:44 Changed 9 months ago by wja

Archiving
rsync to CEDA has failed (script is creating duplicate /tmp/ files).
This error has also been encountered with flight 2010 234e.
Investigating..

comment:45 Changed 9 months ago by wja

Archiving
Previously mentioned error was casued by trailing '/' at end of data path in database.
Currently running rsync to CEDA.

comment:46 Changed 9 months ago by wja

Archiving
RSYNC to CEDA complete (25/02/2019)

Note: See TracTickets for help on using tickets.